OpenStudy (anonymous):

will medal. what is the square root of 96? ^ I I how do you find this out?

Nnesha (nnesha):

factor 96 :-)

OpenStudy (anonymous):

i know 24*4=96 but when i find the sqwr rt. it gives me 9.4756 and so on.

OpenStudy (anonymous):

\[\sqrt{96}= 4\sqrt{6} <--??\]

Nnesha (nnesha):

yep right 96 is not a perfect square so we have to factor it one of number should be perfect square \[\sqrt{24 \times 4 }\] now you can factor 24 too

Nnesha (nnesha):

yep that's right take squareroot o 4 \[2 \sqrt{24} = 2 \sqrt{4 \times 6} \]

OpenStudy (anonymous):

I'm sorry, but this is not clicking in my mind. its 7:40 am here and my mind is still half asleep

OpenStudy (anonymous):

here, ill make a random equation. plz explain how to do this. what is the geometric mean of 8 and 18?

Nnesha (nnesha):

alright divide 96 by2  apply exponent rule which x^m times x^n = x^{m+n} when you multiply same bases you should add their exponents \[\sqrt{2 \times 2 \times 2 \times 2 \times 6} = \sqrt{2^2 \times 2^2 \times 6}\] now you cancel out square with square root